OpenStudy (anonymous):

how would i use my birth day to 1.Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day i will reard medal

OpenStudy (anonymous):

1) Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. General quadratic formula ... y = a(x-h)² + k choose a = 1 because it is easy to work with ... y = (x - 8)² + 15

OpenStudy (anonymous):

do u understand?

OpenStudy (anonymous):

yea i do thanks

OpenStudy (anonymous):

k any more questions

OpenStudy (anonymous):

not right now i dont